OpenStudy (anonymous):

the cancer death rate 100,000 people in the United States can be aproximated by the equation y=-2.2x+216, where x is the number of year since 1990. use this equation to determine in whixh year the approximate cancer death rate will be 150? how can i get this answer?

OpenStudy (anonymous):

y = 150 (target number of deaths) 150 = -2.2 x (year-1990) +216 Subtract 216 from each side -66 = -2.2 x (year -1990) divide both sides by -2.2 30 = year-1990 Add 1990 to each side 2020 = year So in 2020 the cancer deaths will be 150 per year.

OpenStudy (anonymous):

i reallu need help trying to find the answer

OpenStudy (anonymous):

y = -2.2x + 216 --> 2.2x = 216 - y --> x = (216 - y) / 2.2 substituting y = 150 --> x = 66 /2.2 = 30 so in year 2020 (1990 +30) cancer death rate will be 150.

OpenStudy (anonymous):

thank you